# Simulating and investigating various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model

### v2

@@ V Concluding discussion and future work
In this paper, the various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model are investigated. And the results show the nontrivial dependence of the hydrogen bond formation on the parameters. Some analytical results of it are derived as follows
@@
  • In Sec. IV.1, the periodical oscillations of closed system are obtained. Once the dissipations of the non-ideal system is introduced, the oscillations disappear and two final results are obtained: the formation of stable hydrogen bond or the complete breaking of hydrogen bond. Different initial states lead to different results: the system tends to form a stable hydrogen bond when the initial state is $|0\rangle|0\rangle|1\rangle$; the system tends to break the hydrogen bond when the initial state is $|1\rangle|1\rangle|0\rangle$.
@@
  • In Sec. IV.2, the effect of interactions on evolution is studied. When the initial state is $|0\rangle|0\rangle|1\rangle$, the larger the interaction strengths, the harder to form a hydrogen bond, indicating that the interactions hinders the formation of stable hydrogen bond. On the contrary, when the initial state is $|1\rangle|1\rangle|0\rangle$, the larger the interaction strengths, the easier to form a hydrogen bond, indicating that the interactions promotes the formation of stable hydrogen bond. The addition of weak inflows promotes the formation of hydrogen bond, regardless of whether the initial state is $|1\rangle|1\rangle|0\rangle$ or $|0\rangle|0\rangle|1\rangle$.
@@
  • In Sec. IV.3, the effect of dissipations on the hydrogen bond formation is studied. When the initial state is $|0\rangle|0\rangle|1\rangle$, the probability of stable hydrogen bond formation is negatively correlated with $\gamma_{isol}^{out}$ and positively correlated with $\gamma_{bond}^{out}$; however, when the initial state is $|1\rangle|1\rangle|0\rangle$, the probability of stable hydrogen bond formation is negatively correlated with both $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$. $\gamma_{phn}^{out}$ promotes the hydrogen bond formation when the initial state is $|0\rangle|0\rangle|1\rangle$, and hinders it when the initial state is $|1\rangle|1\rangle|0\rangle$.
@@
  • In Sec. IV.4, the effect of inflows on the hydrogen bond formation is studied. Regardless of whether the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$, the probability of stable hydrogen bond formation is positively correlated with $\mu_{isol}$ and negatively correlated with $\mu_{bond}$. $\mu_{phn}$ hinders the formation of hydrogen bond no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$.
@@
  • In Sec. IV.5, the effect of external impulses on the hydrogen bond formation is obtained. It is evident that the more free phonons flow in system from the outside, the lower the probability of stable hydrogen bonds formation, and as the number of phonons increases, the probability approaches a limit. The same conclusion as in the previous subsection is obtained: $\mu_{phn}$ hinders the formation of hydrogen bond in the case of large number of phonons.
[…]
@@
Based on the above conclusions, our work demonstrates that the evolution and hydrogen bond formation can be controlled by selectively choosing system parameters. In particular, they are greatly affected by temperature. Although we only studied the dynamic aspects of simple hydrogen bond model, the results we found will be used as a basis to extend the research to more complex chemical and biological model in the future.
@@ Acknowledgements.
The reported study was funded by China Scholarship Council, project numbers 202308091509, 202308091210 , 202108090327 and 202108090483. The authors acknowledge the center for collective use of ultra-high-performance computing resources (https://www.parallel.ru/) at Lomonosov Moscow State University for providing supercomputer resources that contributed to the research results presented in this paper.
@@ Appendix A Quantum master equation for the target model
$i\hbar\dot{\rho}=[H,\rho]+i\mathcal{L}(\rho)$ (27)
where $H$ is Hamiltonian, $\rho$ is density matrix, $\mathcal{L}=L^{in}+L^{out}$. Then
@@
$+\gamma_{isol}^{in}\left(A_{isol}^{\dagger}\rho A_{isol}-\frac{1}{2}\left(A_{isol}A_{isol}^{\dagger}\rho+\rho A_{isol}A_{isol}^{\dagger}\right)\right)+\gamma_{phn}^{in}\left(A_{phn}^{\dagger}\rho A_{phn}-\frac{1}{2}\left(A_{phnl}A_{phn}^{\dagger}\rho+\rho A_{phn}A_{phn}^{\dagger}\right)\right)$
@@
$+\gamma_{isol}^{out}\left(A_{isol}\rho A_{isol}^{\dagger}-\frac{1}{2}\left(A_{isol}^{\dagger}A_{isol}\rho+\rho A_{isol}^{\dagger}A_{isol}\right)\right)+\gamma_{phn}^{out}\left(A_{phn}\rho A_{phn}^{\dagger}-\frac{1}{2}\left(A_{phn}^{\dagger}A_{phn}\rho+\rho A_{phn}^{\dagger}A_{phn}\right)\right)$
where $\gamma_{bond}^{in}=\mu_{bond}\gamma_{bond}^{out},\ \gamma_{isol}^{in}=\mu_{isol}\gamma_{isol}^{out},\ \gamma_{phn}^{in}=\mu_{phn}\gamma_{phn}^{out}$. Here $\mu_{bond}<1,\ \mu_{isol}<1,\ \mu_{phn}<1$, thus $\gamma_{bond}^{in}<\gamma_{bond}^{out},\ \gamma_{isol}^{in}<\gamma_{isol}^{out},\ \gamma_{phn}^{in}<\gamma_{phn}^{out}$.
@@ Appendix B Estimation of physical quantities
$E_{H-bond}$ $=\frac{E_{bond}}{N_{A}}$ (30)
$=\frac{21\times 10^{3}\ \text{[J/mol}]}{6.022\times 10^{23}\text{[molecules/mol]}}$
[…]
Replacing $E$ in Eq. (23) with $E_{H-bond}$, the characteristic time scale is got as follows
$\tau=\frac{\hbar}{E_{H-bond}}=\frac{\hbar\times N_{A}}{E_{bond}}=\frac{1.0545718\times 10^{-34}\ [J\cdot s]\times 6.022\times 10^{23}\text{[molecules/mol}]}{21\times 10^{3}\ \text{[J/mol]}}\approx 3.175316\times 10^{-15}\ [\rm{s}]$ (31)

### v3

@@ V Concluding discussion and future work
In this paper, the various dynamic aspects of the $\rm{H}_{2}\rm{O}$-related hydrogen bond model are investigated. And the results show the nontrivial dependence of the hydrogen bond formation on the parameters. Some analytical results of it are derived as follows
@@
  • In Sec. IV.1, the periodical oscillations of the closed system are obtained. Once the dissipation of the non-ideal system is introduced, the oscillations disappear, and two final results are obtained: the formation of a stable hydrogen bond or the complete breaking of a hydrogen bond. Different initial states lead to different results: the system tends to form a stable hydrogen bond when the initial state is $|0\rangle|0\rangle|1\rangle$; the system tends to break the hydrogen bond when the initial state is $|1\rangle|1\rangle|0\rangle$.
@@
  • In Sec. IV.2, the effect of interactions on evolution is studied. When the initial state is $|0\rangle|0\rangle|1\rangle$, the larger the interaction strengths, the harder it is to form a hydrogen bond, indicating that the interactions hinder the formation of the stable hydrogen bond. On the contrary, when the initial state is $|1\rangle|1\rangle|0\rangle$, the larger the interaction strengths, the easier it is to form a hydrogen bond, indicating that the interactions promote the formation of the stable hydrogen bond. The addition of weak inflows promotes the formation of the hydrogen bond, regardless of whether the initial state is $|1\rangle|1\rangle|0\rangle$ or $|0\rangle|0\rangle|1\rangle$.
@@
  • In Sec. IV.3, the effect of dissipation on the hydrogen bond formation is studied. When the initial state is $|0\rangle|0\rangle|1\rangle$, the probability of the stable hydrogen bond formation is negatively correlated with $\gamma_{isol}^{out}$ and positively correlated with $\gamma_{bond}^{out}$; however, when the initial state is $|1\rangle|1\rangle|0\rangle$, the probability of the stable hydrogen bond formation is negatively correlated with both $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$. $\gamma_{phn}^{out}$ promotes the hydrogen bond formation when the initial state is $|0\rangle|0\rangle|1\rangle$, and hinders it when the initial state is $|1\rangle|1\rangle|0\rangle$.
@@
  • In Sec. IV.4, the effect of inflows on the hydrogen bond formation is studied. Regardless of whether the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$, the probability of the stable hydrogen bond formation is positively correlated with $\mu_{isol}$ and negatively correlated with $\mu_{bond}$. $\mu_{phn}$ hinders the formation of the hydrogen bond no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$.
@@
  • In Sec. IV.5, the effect of external impulses on the hydrogen bond formation is obtained. It is evident that the more free phonons flow into the system from the outside, the lower the probability of the stable hydrogen bond formation, and as the number of phonons increases, the probability approaches a limit. The same conclusion as in the previous subsection is obtained: $\mu_{phn}$ hinders the formation of the hydrogen bond in the case of a large number of phonons.
  • Comparing Secs. IV.1 $\sim$ IV.5, it is obvious that the effect of temperature on quantum evolution and the hydrogen bond formation is far greater than that of other factors.
@@
Based on the above conclusions, our work demonstrates that the evolution and the hydrogen bond formation can be controlled by selectively choosing system parameters. In particular, they are greatly affected by temperature. Although we only studied the dynamic aspects of the simple hydrogen bond model, the results we found will be used as a basis to extend the research to more complex chemical and biological models in the future.
CRediT authorship contribution statement
J.C. You, R. Chen, and W.S. Li contribute equally to this paper: Investigation (equal), Software (equal), Data curation (equal), Visualization (equal), Formal analysis (equal), Writing — original draft (supporting); H.-H. Miao: Conceptualization (supporting), Methodology (supporting), Validation (lead), Writing — original draft (lead), Writing — review & editing (equal), Project administration (lead), Supervision (supporting); Y.I. Ozhigov: Conceptualization (lead), Methodology (lead), Resources (lead), Writing — review & editing (equal), Supervision (lead).
@@ Acknowledgements.
The reported study was funded by the China Scholarship Council, project numbers 202308091509, 202308091210, 202108090327, and 202108090483. The authors acknowledge the Center for Collective Use of Ultra-high-performance Computing Resources (https://parallel.ru/) at Lomonosov Moscow State University for providing supercomputer resources that contributed to the research results presented in this paper.
@@ Appendix A Quantum master equation for the target model
$i\hbar\dot{\rho}=[H,\rho]+i\mathcal{L}(\rho),$ (25)
where $H$ is the Hamiltonian, $\rho$ is the density matrix, and $\mathcal{L}=L^{in}+L^{out}$. Then
@@
$+\gamma_{isol}^{in}\left(A_{isol}^{\dagger}\rho A_{isol}-\frac{1}{2}\left(A_{isol}A_{isol}^{\dagger}\rho+\rho A_{isol}A_{isol}^{\dagger}\right)\right)+\gamma_{phn}^{in}\left(A_{phn}^{\dagger}\rho A_{phn}-\frac{1}{2}\left(A_{phnl}A_{phn}^{\dagger}\rho+\rho A_{phn}A_{phn}^{\dagger}\right)\right),$
@@
$+\gamma_{isol}^{out}\left(A_{isol}\rho A_{isol}^{\dagger}-\frac{1}{2}\left(A_{isol}^{\dagger}A_{isol}\rho+\rho A_{isol}^{\dagger}A_{isol}\right)\right)+\gamma_{phn}^{out}\left(A_{phn}\rho A_{phn}^{\dagger}-\frac{1}{2}\left(A_{phn}^{\dagger}A_{phn}\rho+\rho A_{phn}^{\dagger}A_{phn}\right)\right),$
where $\gamma_{bond}^{in}=\mu_{bond}\gamma_{bond}^{out},\ \gamma_{isol}^{in}=\mu_{isol}\gamma_{isol}^{out},\ \gamma_{phn}^{in}=\mu_{phn}\gamma_{phn}^{out}$. Here, $\mu_{bond}<1,\ \mu_{isol}<1$, and $\mu_{phn}<1$, thus, $\gamma_{bond}^{in}<\gamma_{bond}^{out},\ \gamma_{isol}^{in}<\gamma_{isol}^{out}$, and $\gamma_{phn}^{in}<\gamma_{phn}^{out}$.
@@ Appendix B Estimation of physical quantities
$E_{H-bond}=\frac{E_{bond}}{N_{A}}=\frac{21\times 10^{3}\ \text{[J/mol}]}{6.022\times 10^{23}\text{[molecules/mol]}}\approx 3.32\times 10^{-20}\ [\rm{J}]=0.217655\ \text{[eV]}.$ (28)
Replacing $E$ in $\tau=\frac{\hbar}{E}$ with $E_{H-bond}$, the characteristic time scale is obtained as follows
$\tau=\frac{\hbar}{E_{H-bond}}=\frac{\hbar\times N_{A}}{E_{bond}}=\frac{1.0545718\times 10^{-34}\ [J\cdot s]\times 6.022\times 10^{23}\text{[molecules/mol}]}{21\times 10^{3}\ \text{[J/mol]}}\approx 3.175316\times 10^{-15}\ [\rm{s}].$ (29)